Circular No. 5769

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 R. Tweedy, Steward Observatory; T. Balonek, Colgate University; and

L. Wells, National Optical Astronomy Observatory and Kitt Peak National Observatory, report: “Several consecutive nights of UBVRI CCD photometry with the Burrell Schmidt telescope at Kitt Peak (cf. IAUC 5746, 5754) indicate that the brightness in U is no longer increasing. This indicates the probable onset of the plateau phase in the light curve. The rate of brightening is decreasing at other colors as well. Magnitudes on Apr. 16.15 UT (using Corwin’s reference stars, IAUC 5742, uncorrected for color) are U = 11.35, B = 11.40, V = 10.88, R = 10.55, I = 10.46 (errors ± 0.01-0.02), with the present rates of increase being 0.00, 0.03, 0.04, 0.05, 0.05 mag/day, respectively.”

Further photoelectric photometry (cf. IAUC 5750, 5755, 5758) by P. N. Appleton and J. J. Eitter, Iowa State University (AE; observations made with the E. W. Fick 0.6-m wide-field CCD system; magnitudes referred to Corwin’s stars 0434 and 0928 from IAUC 5742; rms error typically 0.03 mag); by M. Mikołajewski, K. Rumiński, and B. Wikierski, N. Copernicus University, Toruń (M); and by P. Prugniel, Observatoire de Haute Provence (P): Apr. 4.1 UT, V = 11.81, B–V = +0.45, V–R = +0.31, V–I = +0.57 (AE); 5.1, V = 11.96, B–V = +0.34, V–R = +0.45, V–I = +0.66 (AE); 6.1, V = 11.86, V–R = +0.41 (AE); 10.1, V = 11.37, B–V = +0.50, V–R = +0.33, V–I = +0.44 (AE); 13.1, V = 11.07, B–V = +0.51, V–R = +0.32, V–I = +0.43 (AE); 15.87, V = 10.87 ± 0.02, U–B = –0.01, B–V = +0.54, V–R = +0.51, V–I = +0.40 (M); 16.89, V = 10.86 ± 0.01, U–B = –0.11, B–V = +0.57, V–R = +0.49, V–I = +0.42 (M); 17.07, V = 10.86 ± 0.02, U–B = –0.19, B–V = +0.56, V–R = +0.44, V–I = +0.41 (M); 17.1, V = 10.85, B–V = +0.56, V–I = +0.44 (AE); 18.00, B_{c} = 11.64, B_{c} − R_{c} = +0.93 (P); 19.00, B_{c} = 11.69, B_{c} − R_{c} = +1.02 (P); 19.88, V = 10.87 ± 0.01, U–B = –0.02, B–V = +0.68, V–R = +0.57, V–I = +0.57 (M); 20.00, B_{c} = 11.74, B_{c} − R_{c} = +1.05 (P).

NOVA OPHIUCHI 1993 K. Sekiguchi, South African Astronomical Observatory, forwards the

following provisional photometry of N Oph 1993 obtained on Apr. 19.06 UT by F. van Wyk with the 0.5-m reflector at SAAO Sutherland: V = 9.71, B–V = +0.89, U–B = –0.30, V − R_{c} = +0.92, V − I_{c} = +1.70.

1993 April 20 Daniel W. E. Green